# Identifying Video Game Debugging Bottlenecks: An Industry Perspective

### v1

@@ 1. Introduction
Video games are developed by integrating the work of multidisciplinary teams, such as game programmers, audio engineers, game designers (Gregory, 2018) (Karlsson et al., 2023). During game development, teams are involved in the debugging of various types of bugs varying in categories (e.g., graphical, game logic, and crashes) (Truelove et al., 2021). Presently, all conventional debugging techniques used in traditional software (Decasse and Emde, 1988) are also used to debug video games (Adv, 2023) (Gregory, 2018). However, there is no research available to understand where the bottlenecks are when game developers debug.
@@ 2. Background
Common knowledge is that video games vary in a combination of one or many genres such as Action, Adventure, Role-Playing, Simulation, Strategy and Shooter (Arsenault, 2009), (Apperley, 2006), (Gregory, 2018). However, this difference is mainly due to the game artifacts which produce unique visual aesthetics, interactivity demands from players, game rules, and systems (Apperley, 2006).
2.1. Conventional Debugging Techniques
Currently, traditional debugging techniques span program logging, assertions, breakpoints, profiling, and heap analysis (Decasse and Emde, 1988). Similarly, those conventional debugging techniques are also found to be used when debugging video games Adv (2023) and Gregory (2018). However, video games require an additional set of techniques (e.g.: on-screen console, debug draws, debug camera, cheats and data scrubbing) as described by Gregory (2018). Research into advanced debugging techniques for traditional software has led to advances in fault localizations techniques (Wong et al., 2016) but video game debugging has still lengths to cover (Vanègue et al., 2023).
2.2. Artifacts
The game artifacts used in game development vary from game to game according to McDonald et al. (2021)). These can be categorized according to Gregory (2018) into Game Assets, Front-End, Scripts and Gameplay Code, Artificial Intelligence (AI). More specifically, Game Assets are the 2D/3D models, skeleton hierarchy (rig), sounds, visual effects, animations, narrative scripts, game world, and game objects. Front-End is the graphical user interface (GUI), the head-up display (HUD), and the menus. Scripts and Gameplay Code are the instructions to set game design rules, scripted events in the level, game flow, and subsystems’ logic. Artificial intelligence (AI), automates game logic for non-playable characters.
@@ 3.1. Study 1: Bug Database Extraction
The participating studio’s bug management system had 921 bugs based on a snapshot taken on June 11, 2025 and can be seen on Table 1. Bugs varied in status, with some being new, others being worked on, and finally some closed. These bugs were categorized according to the Truelove et al. (2021) taxonomy of 20 categories of bugs. Bugs were logged by the Quality Assurance (QA) team assigned to the project. All bugs contained fields such as description, summary, assignee, severity, and others important to help the studio prioritize efforts. Throughout the validation of the project, two new categories emerged: Object Behavior and World Traversal. Object Behavior is when two or more objects do not behave as intended when interacting. World Traversal is when regions of the map that should be reachable are not due to world navigation limitations. For severity, a low severity is considered an isolated divergence from the intended system behavior. Medium severity is considered to be a divergence from intended behavior that affects multiple systems. High is considered to have a significant impact on the player’s ability to play the game caused by issues in multiple core systems. Critical severity is when the player cannot play the game. Lastly, the assignment of bugs was done by the QA team with the suspicion of who might be able to fix the bug.
@@ 3.2. Study 2: Survey
This research surveyed 20 game developers who actively created artifacts to be integrated in the game. The survey consisted of seven questions with anonymity being respected. The questions focused on asking about their role, years of experience in the role, frequency of participating in debugging, years of experience in using Unreal Engine, collaboration frequency, and their game debugging setup. Of the 21 surveys sent, 20 participants agreed to share their answers, while 1 refused to share their answers. The survey was sent internally by sharing the survey link and 2 months were given to collect responses. As seen in Figure 2, the majority of our participants have between 4 to over 10 years of experience in their current role and as such their preferred debugging techniques may vary based on the games they’ve debugged in the past. On the other hand, they have about 1-6 years of experience using the Unreal Engine to develop games. As such, the roles studied have a solid understanding of how to develop games but may be new to some of the debugging tools available in the Unreal Engine editor.
@@ 3.3. Study 3: Observations
The observational study involved recording the screen of the game developer while debugging. A total of 4 participants (from the subset of 20) were observed where 3 recorded their screen using the Open Broadcaster Software (OBS) solution and 1 participant requested that their screen be recorded through a video call while sharing the screen. Recordings allowed for microphone usage, but no webcam was required as the focus was to analyze the screen, and if needed, their real-time voice-overs sometimes provided extra details of information. A debugging recording session started when the developer began reading the bug and ended either when interrupted or when they completed their debugging session. These sessions ended when either the game developer identified a fix or when they required another role to fix it.
The recordings were obtained in a locally playable format such as MP4 and were stored on the employer’s computer for safety. These were replayed and analyzed three times in order to first understand the themes, secondly map activities, and thirdly to finalize the theme analysis. The activities were first coded in a spreadsheet by entering the activity code of the debugging event based on the coding book in Table 3. Following, we added for each activity the start time, the end time, the tool used, the debugging technique used, the details of the activity and the reason for moving on to the next one. The total amount of recorded episodes was 32 and for a total of 188.7 minutes focusing on Crashes, Object Behaviors, Object State, and UI.
@@ 3.4. Study 4: Post Interviews
The final study involved interviewing 10 participants (a subset of the 20) about the bottlenecks they faced when debugging, what tool shortcomings they saw, and how they collaborated. The interviews were semi-structured by asking a question and letting participants answer based on past experience or a recent debugging episode. On average, the interviews lasted 12 minutes.
@@ 4.1. Bottlenecks in Video Game Debugging
After analyzing the observations, we identified two main bottlenecks during debugging: Inspection activities totaling 41. 7% and testing activities totaling 35% of the total debugging time. Specifically, the two most time-consuming activities were inspecting components (e.g., game assets), which took 26. 4% of total time, and testing the bug, which took 25. 6% of total debugging time. Figure 4 identifies that collaboration (9% of total time), misc (5.4% of total time), edit (3.7% of total time) and consult (2.3% of total time) were not major bottlenecks when debugging.
Activity Total Time (min) % of Total Sub-Activity Time (min) % of Total
@@
The most time-consuming theme was system localization, taking a total of 108 minutes from the total of 188.7 minutes of debugging time. The second most time-consuming theme was bug understanding, which took 70 minutes. The third was the time taken to fix the fault in 6 minutes. Lastly, 2 minutes of interruptions. In terms of which activities were bottlenecks to these goals, we can see in Figure 1, inspecting components was a common activity (40 min) when trying to localize the system at fault shared with trying to understand a bug. Inversely, testing a bug was more time consuming when trying to understand a bug than when trying to localize the system at fault. We can safely say inspection activities bottleneck system localization but testing ones bottleneck bug understanding.
@@ 4.2. Impact of Debugging Tools and Data
Figure 2. Elapsed Time (%) Spent on Tool by Activity
@@ 4.3. Influence of Cross-Disciplinary Collaboration
From the 10 game developers who were interviewed, we gathered that technical roles were often those involved in debugging. Technical Artists (TAs) fixed issues 50-70% independently, handling data problems while delegating 30% of code issues due to skill gaps. Technical Level Designers (TLDs) had about 50% autonomy. Technical Game Designers (TGDs) exhibited 80% autonomy, delegating 20% after confirming someone else (typically a Programmer) could fix the fault. Game Designers (GDs) were autonomous during the early phases of the project (80%) due to Blueprint interactions, dropping to 60% in later stages requiring the help of the programmer to complete the logic implementation. Level Designers (LDs) fixed 40% of issues, while Artists (A) resolved almost 100% of their bugs independently when it came to Art asset changes in a DCC tool. Gameplay Programmers (GPPs) maintained 90% autonomy, delegating 10% of data issues that game designers used to configure systems. Engine Programmers (EPs) often support other teams that need engine work and are 100% autonomous when they need to fix their own systems.
Figure 3 displays the collaboration ratios of the roles when debugging. By analyzing the debugging sessions and interviews, Gameplay Programmers, Engine Programmers and Level Designers were at the core of collaboration requests. Game Designers engaged frequently with Gameplay Programmers (e.g. approximately 20-25% of their collaborations) to resolve gameplay logic bugs. Meanwhile, Level Designers worked with Technical Level Designers (e.g. around 10%) for level-related issues. Technical Artists often collaborated with Artists, (e.g., 20%) to support asset integration and rendering matters. Technical Artists also collaborated with Gameplay Programmers (e.g. 10%), when the gameplay logic was linked to their asset integration. Engine Programmers directly collaborated with Gameplay Programmers due to the Engine Programmers creating the underlying systems being used by the gameplay logic.
Figure 3. Cross-Collaboration Frequency Ratios
@@ 5. Threats to Validity
Our research project is based on the analysis of a video game in development. As such, the focus is mostly on the gameplay aspect and relies significantly on game designers, gameplay programmers, engine programmers, and technical designers. Consequently, other roles such as technical artists, level designers, animators, and sound engineers do not debug as much as the previous roles. Consequently, there is a bias on the types of debugging sessions that were observed which focused on object related bugs, crashes and HUD issues. In subsequent phases, the focus will be on debugging performance degradations, level inconsistencies, ai, etc.
@@
Lastly, this research involves participants that are considered veterans. Replicating the studies may prove difficult and results may therefore vary in regards to time to debug certain bugs. In this paper, the participants understood the debugging techniques to use, and consequently, we did not extract data on their application. Instead, our focus is on identifying bottlenecks during the entire debugging process. Similarly, the lack of data on how developers fix bugs is not part of our study, since most often fixing was considered as a trivial aspect of debugging for developers. As well, collaboration is key in this studio since different roles work on different aspects of the game and knowledge of a system is therefore spread out.
@@ 6. Discussions
Developers that debug are often doing so from the Engine editor’s tools. In some cases, they integrate their editor with external debugging tools by attaching it to the game process. These tools require technical knowledge and can bottleneck the debugging of bugs if experience is lacking. Likewise, the data extracted during these debugging sessions are currently not structured to facilitate debugging. It is not uncommon for developers to extend the engine to support pain points faced during debugging. This could lead to ”rubberband” solutions in which developers may talk to the data to help them better understand the systems.
@@ 7. Related Work
Compared to Alaboudi and LaToza (2023), the bottlenecks that affect most traditional software and video games are the debugging time spent on testing activities with 20% and 35%, respectively. Politowski et al. (2021) studied 200 postmortems, uncovering that 46% of the problems faced during game development lies in the use of tools. Many authors have investigated how to improve the tooling for game development. For example, Machado et al. (2019) set out to improve game design prototyping by using an AI that recommends existing game mechanics. Lucas and Martinho (2017), proposed 3Buddy, a co-creative level design tool aimed at assisting game developers in generating innovative and engaging levels. Engström (2019) proposed Deig Writing Companion that enables script iterations and writing dialogues based on play-through traces, similar to a writing a screenplay. In addition to that, there are external tools that already support game development, such as Wwise for audio (Wang and Olivieri, 2018) and Maya and 3DS Max for animation (Cooper, 2021). However, according to the latest study on debugging tools, Vanègue et al. (2023) did not identify any tools that support bug reproduction activities.

### v2

@@ 1. Introduction
Video games are developed by integrating the work of multidisciplinary teams, such as game programmers, audio engineers, game designers (Gregory, 2018)(Karlsson et al., 2023). During game development, teams are involved in the debugging of various types of bugs varying in categories (e.g., graphical, game logic, and crashes) (Truelove et al., 2021). Presently, all conventional debugging techniques used in traditional software (Decasse and Emde, 1988) are also used to debug video games (1)(Gregory, 2018). However, there is no research available to understand where the bottlenecks are when game developers debug.
@@
In this paper, we find it relevant to first support its content with a brief background on the development of video games. Following, we plan to describe the methods used to retrieve the necessary data to evaluate our three research questions. Specifically, we will present our mixed research approach involving a bug database analysis and three studies on participants such as a survey, observation of debugging sessions, and lastly a semiformal interview. We will then present the results obtained from performing these studies. The following section will be on threats to validity since these results pertain to a specific studio and population. Lastly, we will discuss these industry observations and the conclusions that can be drawn from them.
@@ 2. Background
Common knowledge is that video games vary in a combination of one or many genres such as Action, Adventure, Role-Playing, Simulation, Strategy and Shooter (Arsenault, 2009)(Apperley, 2006)(Gregory, 2018). However, this difference is mainly due to the game artifacts which produce unique visual aesthetics, interactivity demands from players, game rules, and systems (Apperley, 2006).
The game artifacts used in game development vary from game to game according to McDonald et al. (2021). These can be categorized into Game Assets, Front-End, Scripts and Gameplay Code, Artificial Intelligence (AI) as stated by Gregory (2018). More specifically, Game Assets are the 2D/3D models, skeleton hierarchy (rig), sounds, visual effects, animations, narrative scripts, game world, and game objects. Front-End is the graphical user interface (GUI), the head-up display (HUD), and the menus. Scripts and Gameplay Code are the instructions to set game design rules, scripted events in the level, game flow, and subsystems’ logic. Artificial intelligence (AI), automates game logic for non-playable characters.
Currently, traditional debugging techniques span program logging, assertions, breakpoints, profiling, and heap analysis (Decasse and Emde, 1988). Similarly, these conventional debugging techniques are also used when debugging video games (1)(Gregory, 2018). However, video games require an additional set of techniques (e.g.: on-screen console, debug draws, debug camera, cheats and data scrubbing) as described by Gregory (2018). Research into advanced debugging techniques for traditional software has led to advances in fault localizations techniques (Wong et al., 2016) but video game debugging has still lengths to cover (Vanègue et al., 2023).
@@ 3.1. Study 1: Bug Database Extraction
The participating studio’s bug management system had 921 bugs based on a snapshot taken on June 11, 2025 and can be seen on Table 1 on page 1. Bugs varied in status, with some being new, others being worked on, and finally some closed. These bugs were categorized according to the taxonomy of 20 categories of bugs provided by Truelove et al. (2021). Bugs were logged by the Quality Assurance (QA) team assigned to the project. All bugs contained fields such as description, summary, assignee, severity, and others important to help the studio prioritize efforts. Throughout the validation of the project, two new categories emerged: Object Behavior and World Traversal. Object Behavior is when two or more objects do not behave as intended when interacting. World Traversal is when regions of the map that should be reachable are not due to world navigation limitations. For severity, a low severity is considered an isolated divergence from the intended system behavior. Medium severity is considered to be a divergence from intended behavior that affects multiple systems. High is considered to have a significant impact on the player’s ability to play the game caused by issues in multiple core systems. Critical severity is when the player cannot play the game. Lastly, the assignment of bugs was done by the QA team with the suspicion of who might be able to fix the bug.
@@
A table with Bug Categories and respective occurrences count by severity. The last column is the Total count.
@@ 3.2. Study 2: Survey
This research surveyed 20 game developers who actively created artifacts to be integrated in the game. This narrows down the participation to seven roles having a technical background or those with the ability to submit changes to the engine through tools. The survey consisted of three questions with anonymity being respected. The questions focused on asking about their role, years of experience in the role, and years of experience in using Unreal Engine. Of the 21 surveys sent, 20 participants agreed to share their answers, while 1 refused to share their answers. The survey was sent internally by sharing the survey link and 2 months were given to collect responses.
As seen in Figure 2 on page 2, the majority of our participants have between 4 to over 10 years of experience in their current role and as such their preferred debugging techniques may vary based on the games they’ve debugged in the past. On the other hand, they have about 1-6 years of experience using the Unreal Engine to develop games. As such, the roles studied have a solid understanding of how to develop games but may be new to some of the debugging tools available in the Unreal Engine editor.
@@
A table withe the list of Participants from this organized by ID, Role, Years of Experience in Role and Years of Experience using the Unreal Engine.
@@ 3.3. Study 3: Observations
The observational study involved recording the screen of the game developer while debugging. This study focused on the most frequent and severe bugs encountered during development, since those are the ones being debugged actively and had the highest chance of being fixed during the time provided for this study. As such, not at all participants were involved in the debugging of those categories. In this case, only a total of 4 participants (from the subset of 20) participated in having their screens recorded while they debugged. 3 of those participants recorded their screen using the Open Broadcaster Software (OBS) solution and 1 participant requested that their screen be recorded through a video call while sharing the screen. Recordings allowed for microphone usage, but no webcam was required as the focus was to analyze the screen, and if needed, their real-time voice-overs sometimes provided extra details of information. A debugging recording session started when the developer began reading the bug and ended either when interrupted or when they completed their debugging session. These sessions ended when either the game developer identified a fix or when they required another role to fix it.
The recordings were obtained in a locally playable format such as MP4 and were stored on the employer’s computer for safety. These were replayed and analyzed three times in order to first understand the themes, secondly map activities, and thirdly to finalize the theme analysis. The activities were first coded in a spreadsheet by entering the activity code of the debugging event based on the coding book in Table 3 on page 3. Following, we added for each activity the start time, the end time, the tool used, the debugging technique used, the details of the activity and the reason for moving on to the next one. The total amount of recorded episodes was 32 and for a total of 188.7 minutes focusing on Crashes, Object Behaviors, Object State, and UI.
@@
A table describing the Coding Book used to thematically organize the activities observed during the Debugging Sessions.
@@ 3.4. Study 4: Post Interviews
The final study involved interviewing at least one participant from each of the seven roles identified previously. For this study, 10 participants (a subset of the 20) wished to be interviewed about the bottlenecks they faced when debugging, what tool shortcomings they saw, and how they collaborated. The interviews were semi-structured by asking a question and letting participants answer based on past experience or a recent debugging episode. On average, the interviews lasted 12 minutes.
@@ 4.1. Bottlenecks in Video Game Debugging
After analyzing the observations, we identified two main bottlenecks during debugging: Inspection activities totaling 41.7% and testing activities totaling 35% of the total debugging time. Specifically, the two most time-consuming activities were inspecting components (e.g., game assets), which took 26.4% of total time, and testing the bug, which took 25.6% of total debugging time. Figure 4 on page 4 identifies that collaboration (9% of total time), misc (5.4% of total time), edit (3.7% of total time) and consult (2.3% of total time) were not major bottlenecks when debugging.
Activity Total Time (min) % of Total Sub-Activity Total Time (min) % of Total
@@
A table describing the time the participants spent in each type of activities. On the left half, you have the activity types with its total time and percentage. On the right half you have the sub-activities with the total time and percentage of total as well.
The most time-consuming theme was system localization, taking a total of 108 minutes from the total of 188.7 minutes of debugging time. The second most time-consuming theme was bug understanding, which took 70 minutes. The third was the time taken to fix the fault in 6 minutes. Lastly, 2 minutes of interruptions. In terms of which activities were bottlenecks to these goals, we can see in Figure 1 on page 1, inspecting components was a common activity (40 min) when trying to localize the system at fault shared with trying to understand a bug. Inversely, testing a bug was more time consuming when trying to understand a bug than when trying to localize the system at fault. We can safely say inspection activities bottleneck system localization but testing ones bottleneck bug understanding.
@@
A heatmap with the yellow and green colors identifying the activities (y-axis) with the most time spent in relation to the themes (x-axis) identified in the coding book.
Figure 2. Elapsed Time (%) Spent on Tool by Activity
A heatmap with the yellow and green colors identifying the activities (y-axis) with the most time spent in relation to the tools being used (x-axis).
Figure 3. Cross-Collaboration Frequency Ratios
Directed Graph with the ratio of collaboration between roles from our study.
@@ 4.3. Influence of Cross-Disciplinary Collaboration
From the 10 game developers who were interviewed, we gathered that technical roles were often those involved in debugging. Technical Artists (TAs) fixed issues 50-70% independently, handling data problems while delegating 30% of code issues due to skill gaps. Technical Level Designers (TLDs) had about 50% autonomy. Technical Game Designers (TGDs) exhibited 80% autonomy, delegating 20% after confirming someone else (typically a Programmer) could fix the fault. Game Designers (GDs) were 80% of the time autonomous during the early phases of the project due to Blueprint interactions, dropping to 60% in later stages requiring the help of the programmer to complete the logic implementation. Level Designers (LDs) fixed 40% of issues, while Artists (A) resolved almost 100% of their bugs independently when it came to Art asset changes in a DCC tool. Gameplay Programmers (GPPs) maintained 90% autonomy, delegating 10% of data issues that game designers used to configure systems. Engine Programmers (EPs) often support other teams that need engine work and are 100% autonomous when they need to fix their own systems.
Figure 3 on page 3 displays the collaboration ratios of the roles when debugging. By analyzing the debugging sessions and interviews, Gameplay Programmers, Engine Programmers and Level Designers were at the core of collaboration requests. Game Designers engaged frequently with Gameplay Programmers (e.g. approximately 20-25% of their collaborations) to resolve gameplay logic bugs. Meanwhile, Level Designers worked with Technical Level Designers (e.g. around 10%) for level-related issues. Technical Artists often collaborated with Artists, (e.g., 20%) to support asset integration and rendering matters. Technical Artists also collaborated with Gameplay Programmers (e.g. 10%), when the gameplay logic was linked to their asset integration. Engine Programmers directly collaborated with Gameplay Programmers due to the Engine Programmers creating the underlying systems being used by the gameplay logic.
@@ 5. Threats to Validity
Our research project is based on the analysis of a video game in development. As such, the focus is mostly on the gameplay aspect and relies significantly on game designers, gameplay programmers, engine programmers, and technical designers. Consequently, other roles such as technical artists, level designers, animators, and sound engineers were not studied proportionally to the previous roles. Consequently, there is a bias on the types of debugging sessions that were observed which focused on object related bugs, crashes, and HUD issues. In subsequent phases, the focus will be on debugging performance degradations, level inconsistencies, ai, etc.
@@
Lastly, this research involves participants who are veterans in the industry and currently work on the same project within the same studio. As such, the findings may be difficult to generalize due to the consequent small population. Nevertheless, in this paper, the evaluation of the debugging sessions through our coding book is applicable to other video game debugging sessions notwithstanding the studio or game genre. Furthermore, the fact that the Misc, Edit, and Consult activity categories are small in proportion to other categories means that a more diverse population is needed.
@@ 6. Discussions
Developers that debug are often doing so from the Engine editor’s tools. In some cases, they integrate their editor with external debugging tools by attaching it to the game process. These tools require technical knowledge and can bottleneck the debugging of bugs if experience is lacking. Likewise, the data extracted during these debugging sessions are currently not structured to facilitate debugging. It is not uncommon for developers to extend the engine to support pain points faced during debugging. This could lead to ”rubberband” solutions in which developers may talk to the data to help them better understand the systems. Furthermore, these findings could also be extended to other game engines that expose their systems through an editor such as Unity and Godot.
@@ 7. Related Work
Compared to Alaboudi and LaToza (2023), the bottlenecks that affect most traditional software and video games are the debugging time spent on testing activities with 20% and 35%, respectively. Politowski et al. (2021) studied 200 postmortems, uncovering that 46% of the problems faced during game development lies in the use of tools. Many authors have investigated how to improve the tooling for game development. For example, Machado et al. (2019) set out to improve game design prototyping by using an AI that recommends existing game mechanics. Lucas and Martinho (2017) proposed 3Buddy, a co-creative level design tool aimed at assisting game developers in generating innovative and engaging levels. Engström (2019) proposed Deig Writing Companion that enables script iterations and writing dialogues based on play-through traces, similar to a writing a screenplay. In addition to that, there are external tools that already support game development, such as Wwise for audio (Wang and Olivieri, 2018) and Maya and 3DS Max for animation (Cooper, 2021). However, according to the latest study on debugging tools, Vanègue et al. (2023) did not identify any tools that support bug reproduction activities.